### v1

# SounDiT: Geo-Contextual Soundscape-to-Landscape Generation

Abstract
We present a novel and practically significant problem—Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation—which aims to synthesize geographically realistic landscape images from environmental soundscapes. Prior audio-to-image generation methods typically rely on general-purpose datasets and overlook geographic and environmental contexts, resulting in unrealistic images that are misaligned with real-world environmental settings. To address this limitation, we introduce a novel geo-contextual computational framework that explicitly integrates geographic knowledge into multimodal generative modeling. We construct two large-scale geo-contextual multimodal datasets, SoundingSVI and SonicUrban, pairing diverse soundscapes with real-world landscape images. We propose SounDiT, a novel Diffusion Transformer (DiT)-based model that incorporates geo-contextual scene conditioning to synthesize geographically coherent landscape images. Furthermore, we propose a practically-informed geo-contextual evaluation framework, the Place Similarity Score (PSS), across element-, scene-, and human perception-levels to measure consistency between input soundscapes and generated landscape images. Extensive experiments demonstrate that SounDiT outperforms existing baselines in both visual fidelity and geographic settings. Our work not only establishes foundational benchmarks for GeoS2L generation but also highlights the importance of incorporating geographic domain knowledge in advancing multimodal generative models, opening new directions at the intersection of generative AI, geography, urban planning, and environmental sciences.
Figure 1: Geo-contextual soundscape-to-landscape (GeoS2L) generation aims to synthesize realistic landscape images from environmental soundscapes. We introduce large-scale geo-contextual datasets, the SounDiT model, and the Place Similarity Score evaluation framework to support this task.
1 Introduction
“The earth has music for those who listen.” – William Shakespeare
Each place is characterized by a unique set of visual and sounding environments [66, 4, 6, 2], such as natural bird calls in a quiet park starkly contrasting with bustling traffic in downtowns at peak hours. Understanding such geo-contextual linkages regarding how visual and sounding environments co-occur benefits broad applications in geography [74, 14, 39], environmental psychology [53, 26], ecology [60], and urban planning [15]. Integrating these modalities not only advances our knowledge of human-environment interactions [76], but also supports real-world practices [33]. For instance, understanding how the environment shapes human multi-sensory experiences and perceptions [75, 86], informing noise mitigation and place-making practices to benefit human well-being [1, 32], and enhancing inclusive navigation for visually impaired populations [37, 8].
Prior research in the Audio-to-Image (audio2image) stream that translates auditory information into visual representations has offered promising opportunities [70, 61, 80, 79, 5]. However, we observed two noticeable gaps when applying prior studies in geographic contexts and planning practices. First, most audio2image studies rely on generic audio-visual datasets such as object audio [57, 85, 9], human voice [54, 49], weathers [42], and limited scene types [41], and overlook geo-contextual insights such as parks vs. beaches and urban vs. rural settings [81] (See Appendix LABEL:appendix:dataset). Geographers are primarily interested in the environmental context of a soundscape rather than identifying sound sources such as objects or animals. Second, common evaluation metrics employed in prior studies, such as the Fréchet Inception Distance (FID) [30], may fall short in capturing geographic semantic alignment between the input soundscape and generated images. Ideally, the evaluation metrics should consider environmental characteristics and geographic contexts: the place setting depicted in the input soundscape and those represented in the generated images should be similar to ensure geographic contextual coherence and relevance. Furthermore, Diffusion Transformer (DiT) [59] techniques have recently demonstrated impressive results across various tasks such as image [44, 10, 11] and video generations [51, 77, 38, 48]. However, they remain unexplored in audio2image generation and have yet to incorporate geographic knowledge to support applications in geography and urban planning.
Given these challenges, we introduce and formalize a new problem, namely, Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation (Figure 1), to synthesize realistic landscape images from environmental soundscapes, optionally conditioned on scene context. The GeoS2L problem extends traditional audio2image synthesis by incorporating geo-contextual insights with significant practical implications. Here, landscape refers to geographical environments shaped by natural and built features [52]. A soundscape denotes the acoustic environment as perceived by humans at places, comprising sounds from both natural and anthropogenic sources (e.g., bird calls and vehicle noises) [68]. A scene indicates place types such as beach, forest, street, to guide image generation. Our contributions span three key areas: First, we introduce two large-scale, multi-modal geo-contextual datasets: SoundingSVI and SonicUrban, composed of over 330K and 250K soundscape-landscape image pairs. These datasets cover diverse geographic locations and scenes, supporting geo-contextual realistic synthesis. Second, we develop an innovative generative model, SounDiT, which integrates explicit geographic contexts into DiT to effectively translate environmental soundscapes into corresponding geographically coherent landscape images. Third, we design a novel geo-contextual evaluation framework, namely, Place Similarity Scores (PSS), which includes three metrics: element-level, scene-level, and perception-level similarity to assess geographic contexts and environmental consistency of generated landscape images. Our proposed SounDiT achieves state-of-the-art performance on SoundingSVI and SonicUrban. Beyond technical contributions, this work supports applications in geography, environment, and urban planning with significant practical impacts.
2 Related Work
Audio-to-Image Synthesis aims to translate acoustic features into coherent and meaningful images and has gained increasing attention with the advancement of various multi-modal alignment foundation models [79, 70] and generative techniques [24, 65]. Current audio2image models are broadly divided into GAN-based (e.g., sound2scene [70]) and Diffusion-based (e.g., AudioToken [80], and GluenGen [61]). These models mainly rely on generic datasets which lack geo-contextual information, often limited to object audios [57], human voice [54], weather [42], and broad scene types [41]. As a result, they often produce stylized or unrealistic outputs and fail to capture the geographic coherence between soundscapes and landscapes (See Appendix LABEL:appendix:dataset).
Landscapes and Soundscapes represent two dimensions for characterizing places [46, 19]. Landscapes offer visual representations of the physical layout and built environments, and are commonly captured via street view imagery (SVI) and remote sensing (RS) images [41, 42]. In parallel, soundscapes refer to the acoustic environment perceived at specific places. Existing datasets, such as BirdCLEF [34], BirdSet [64], SoundingEarth [28]), AudioSet [20], UrbanSound8K [67], SONYC [7], EnigenScape [25], and EMO-Soundscape [16], have been utilized to support tasks including monitoring biodiversity [63], noise pollution [32], and human perception analysis [18]. However, prior audio2image studies have overlooked such geographic and environmental insights, limiting their relevance for real-world applications. To address this gap, our work integrates geo-contextual information for multimodal generative modeling by constructing datasets and evaluation frameworks.
3 Method
3.1 Preliminary
Diffusion Model. We adopt a latent-based diffusion model, consisting of a forward diffusion and backward denoising process conducted in the latent space $\mathbf{z}$. The forward diffusion $\mathbf{q}$ starts from the initial latent $\mathbf{z}_{0}$, gradually adding Gaussian noise over timestep $T$ based on a fixed variance $\alpha$ and defined as:
$q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1})=\mathcal{N}\bigl{(}\mathbf{z}_{t};\sqrt{\alpha_{t}}\,\mathbf{z}_{t-1},(1-\alpha_{t})\,\mathbf{I}\bigr{)},\quad t=1,\dots,T.$
The backward denoising process $p_{\theta}$ is trained to predict the less noisy latent:
$p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})=\mathcal{N}\bigl{(}\mathbf{z}_{t-1};\,\mu_{\theta}(\mathbf{z}_{t},t,v),\,\Sigma_{\theta}(\mathbf{z}_{t},t,v)\bigr{)},$
where $v$ is the conditional signal and $\mu_{\theta}$, $\Sigma_{\theta}$ are predicted through a denoising network $\epsilon_{\theta}(\mathbf{z}_{t},t,v)$.
Classifier-Free Guidance. To control the influence of the conditioning signal $v$ during backward denoising, we employ classifier-free guidance (CFG) to predict noise $\hat{\boldsymbol{\epsilon}}$ based on both conditional and unconditional signals with specific guidance scale $s$.
$\hat{\boldsymbol{\epsilon}}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)+s\bigl{[}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,v)-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)\bigr{]}.$
3.2 Geo-Contextual Soundscape-to-Landscape Generation (GeoS2L) Problem Definition
In this work, we formulate our task as the Geo-Contextual Soundscape-to-Landscape (GeoS2L) problem, a specialized variant of the audio2image problem, which aims to generate landscape images from environmental soundscapes guided by an optional scene prompt. We define a landscape $l$ as the geographical and ecological environment shaped by natural and built features, and a soundscape $s$ as the human-perceived acoustic environment at a given place, comprising both natural and anthropogenic sounds such as bird calls and traffic noise. To better support real-world practices, we introduce scene prompt $c$ as an optional semantic label (e.g., park, beach, street) to provide additional geographic semantic context.
Given a dataset $D=\{\,(s_{i},c_{i},l_{i})\mid i=1,\dots,N\},$ our goal is to learn a generator $G\colon(s_{i},c_{i})\mapsto\hat{l}_{i}$ such that the synthesized image $\hat{l}_{i}$ is both visually realistic and geo-contextually consistent to the ground truth $l_{i}$, quantified by relevance function $R(s,c,l)$. This can be formulated as:
$\mathcal{L}\;=\;\mathbb{E}_{(s_{i},c_{i},l_{i})\sim D}\Bigl{[}\,R(s_{i},c_{i},l_{i})\;-\;R\bigl{(}s_{i},c_{i},\hat{l}_{i}\bigr{)}\Bigr{]}.$
Minimizing $\mathcal{L}$ enables the model to capture cross-modal correlations among sounding, semantic and visual features, thereby producing landscape images that reflect both the acoustic characteristics of soundscape and intended scene.
Figure 2: Framework of SounDiT. To effectively capture geographic contexts, SounDiT integrates soundscape, scene, and landscape embeddings from pre-trained encoders as inputs. The landscape embedding is fused with the timestep embedding and fed into a self-attention layer to extract visual features. The scene embedding then guides conditioning via a cross-attention layer. The TA-PMoE module further combines these features with soundscape features. During inference, a geo-contextual RAG module provides additional geographic contexts to enhance alignment with real-world environmental settings.
3.3 SounDiT
To support the GeoS2L task, we propose SounDiT, a diffusion-based generative model that synthesizes geographically coherent landscape images conditioned on environmental soundscapes and optional scene prompts (Figure 2). SounDiT advances geo-contextual landscape generation through two key innovations: (1) the SounDiT Block, which integrates sound, scene, and visual cues into the denoising process; and (2) a geo-contextual Retrieval-Augmented Generation (RAG) to guide image generation with additional geographic contexts at inference time.
Multimodal Encoders. To incorporate acoustic, semantic and visual features into SounDiT, we employ three encoder modules: (1) a Soundscape Encoder $\mathcal{E}_{\mathrm{Sound}}$, (2) a Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ and (3) a Landscape Encoder–Decoder pair $\bigl{(}\mathcal{E}_{L},\mathcal{E}^{\prime}_{L}\bigr{)}$. The Soundscape Encoder $\mathcal{E}_{Sound}$ embeds soundscape $s$ into an audio embedding $e_{\mathrm{sound}}$. The Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ encodes the optional scene prompt $c$ into a text embedding $e_{\mathrm{scene}}$. The Landscape Encoder $\mathcal{E}_{L}$ compresses a landscape image $l_{i}\in\mathbb{R}^{H\times W\times 3}$ into a latent embedding $e_{\mathrm{landscape}}$, and the Landscape Decoder $\mathcal{E^{\prime}}_{L}$ reconstructs the latent embedding back into a landscape image $\hat{l_{i}}\in\mathbb{R}^{H\times W\times 3}$.
SounDiT Block. Despite the demonstrated effectiveness of DiT in image generation, they have not been integrated into audio2image tasks. To effectively model geographic contexts from both soundscape and scene inputs, we propose the SounDiT block, a geo-contextual DiT to generate landscape images that align with intended geographic settings. The SounDiT operates in three stages: First, the timestep embedding $e_{\mathrm{t}}$ is fused with landscape latent embedding $e_{\mathrm{landscape}}$, and fed into a multi-head self-attention layer to capture visual features. Second, a multi-head cross-attention conditioned on the scene embedding $e_{\mathrm{scene}}$ is applied to enrich these features with scene semantics, producing scene-landscape representation. Third, to further integrate multi-level soundscape features, we introduce Time-Aware Prototype Mixture-of-Experts (TA-PMoE) module, composed of $M$ parallel cross-attention layers as experts. Each cross-attention layer fuses the scene-landscape representation with the soundscape embedding $e_{\mathrm{sound}}$, resulting in $M$ soundscape-scene-landscape representations $\mathbf{P}=[\,\mathbf{p}_{1},\dots,\mathbf{p}_{M}\,]$. A router $\mathbf{w}$ then calculates a weighted combination of these representations $\mathbf{P}$ based on temperature-scaled Softmax:
[…]
Here, $\tau>0$ controls the sharpness of the distribution: smaller values of $\tau$ concentrate weight on the most relevant soundscape-scene-landscape representation, while larger values lead to a smoother allocation across representations. Finally, the combined soundscape-scene-landscape representation is passed through a pointwise feedforward network to produce the denoised output. By hierarchically fusing visual, scene, and soundscape representations, SoundDiT effectively generates geo-contextually coherent landscape images.
Geo-contextual RAG Inference. In practice, additional geo-contextual information, such as location, scene and spatial layout [47], may be collected alongside soundscapes and landscapes to guide image generation. Thus, we propose a flexible geo-contextual RAG pipeline to further enhance the geographic coherence of the generated landscape images. Our method consists of two stages: (1) Geo-contextual Knowledge Database (GeoKD), where we store soundscape embeddings along with associated scene categories and layout representations with geographic locations; and (2) Geo-contextual Initialization, where the most relevant scene and layout are retrieved based on soundscape similarity and geographic proximity. These retrieved cues are then used to initialize the diffusion generation to enhance geographic coherence.
4 Datasets: SoundingSVI and SonicUrban
Prior audio2image studies have primarily relied on general datasets such as VEGAS [85], VGGSound [9], and ACAV100M [40]. Despite their proven effectiveness, these datasets lack geographic contexts. For instance, geographers may be more interested in generating a wetland implied by a frog call, rather than an image of the frog itself. A car horn might suggest a congested urban street, rather than a static vehicle. Some datasets, such as Landscape [41] and IntoTheWild [42], have included paired soundscape and image data, but offer limited spatial coverage and scene diversity. To address this gap, we introduce two new multimodal datasets: SoundingSVI and SonicUrban, designed to support GeoS2L generation with large-scale, geographically diverse, and context-rich data. Figure 3 provides several sample images of these two datasets.
[…]
Dataset Modalities # Size Geo-Contextual Scene Diversity Spatial Coverage Viewpoint
ACAV100M [40] Video 100M -- -- -- Ground-level
VEGAS [85] Video 28K -- -- -- Ground-level
[…]
IntoTheWild [42] Video 142 $\checkmark$ Low -- Ground-level
Landscape [41] Video 1747 $\checkmark$ Medium -- Ground-level
SoundingEarth [28] Audio & Image 50K $\checkmark$ High High Overhead
SoundingSVI (Ours) Audio & Image 330K $\checkmark$ High High Ground-level
SonicUrban (Ours) Video 250K $\checkmark$ High High Ground-level
Figure 3: Representative samples from the SoundingSVI and SonicUrban datasets, showing diverse geographic contexts.
SoundingSVI. The SoundingSVI dataset was constructed from geotagged soundscape recordings from the Aporee platform [56]. For each raw soundscape recording, we retrieved nearby Google Street View Image (SVI) using its latitude and longitude to capture the surrounding visual landscapes. This follows a similar approach to the Sounding Earth dataset [28], which paired geotagged audio with remote sensing imagery. The raw recordings were then segmented into 10-second clips, with a human voice detection model [73] used to filter clips dominated by human voice. Notably, multiple candidate SVIs captured from different angles and timestamps might be retrieved for clips from the same recording. A sound source localization model [58] was further applied to match each clip with the most relevant landscape image. In total, SoundingSVI dataset comprises 334,495 soundscape-landscape pairs in 90 countries. To support optional scene prompts during training, a scene type for each soundscape-landscape pair was predicted by a scene classification model [27].
SonicUrban. The SonicUrban dataset leverages unedited videos as a source of temporally and spatially consistent soundscape-landscape pairs. Following a prior practice [86], we searched and manually checked YouTube videos that contain real-world soundscapes and landscapes, using a combination of city names and keywords, like “City walk; New York”. Each video was processed into 10-second audio clips, and 10 evenly distributed frames were extracted for each clip. The same human voice detection model [73] and sound source localization model [58] used in SoundingSVI were applied to filter human-voice dominated clips and identify the most representative frame for each audio clip, respectively. In total, the constructed SonicUrban consists of 250,310 soundscape-landscape pairs across 131 cities and 97 countries. A scene classification model [27] was further employed to generate a scene prompt for each pair as a geographic scene condition.
5 Geo-Contextual Evaluation Framework
Figure 4: Geo-contextual evaluation framework. The ground truth and output landscape images are assessed at the element-, scene-, and human-perception levels following geography and urban planning practices.
Prior audio2image studies commonly rely on evaluation metrics such as Fréchet Inception Distance (FID) [30], Audio-Image Similarity (AIS) [30], and Image-Image Similarity (IIS) [5], to assess visual or audio fidelity, overlooking the underlying geographic settings and target scenes. Since soundscapes and landscapes co-occur in the same space, they are expected to share similar environmental characteristics and place settings. Following this hypothesis, we introduce a novel practically-informed geo-contextual evaluation framework — Place Similarity Score (PSS) — that measures the environmental settings of soundscape and landscape across element, scene, and human perception levels (Figure 4). These three levels of measurements of environmental characteristics have been widely employed in geographic and urban planning practices [50, 36].
Table 2: Place Similarity Score (PSS)
Score Definition
$\text{PSS}_{\text{element}}$ $\frac{1}{n}\sum_{i=1}^{n}\frac{E_{i}\cdot\hat{E}_{i}}{\|E_{i}\|\|\hat{E}_{i}\|}$
$\text{PSS}_{\text{scene}}$ $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{P_{i}^{\text{k}}\cap T_{i}^{\text{k}}\neq\varnothing\right\}$
$\text{PSS}_{\text{perception}}$ $\frac{1}{n}\sum_{i=1}^{n}\left\|R(l_{i})-R(\hat{l}_{i})\right\|_{1}$
Element-level Evaluation quantifies the similarity of geographic elements (e.g., trees [43], sky [23], water bodies [69], traffic signs [3], and buildings [13]) between the geographical contexts of the input soundscape and generated landscapes [35]. We employ the DeepLabV3 semantic segmentation model [12], pre-trained on the ADE20K dataset [84], to identify 150 predefined geographic elements $E_{i}$ in each image $i$. For each image $i$, we calculate a normalized element distribution vector $E_{i}=\{e_{i}^{1},e_{i}^{2},\dots,e_{i}^{k}\}$, where each $e_{i}^{k}$ represents the ratio of pixels identified as the $k$-th element. [17]. We calculate element ratio vectors $E_{i}$ and $\hat{E_{i}}$ from the ground truth landscape image $l_{i}$ and their corresponding image $\hat{l_{i}}$. We then measure the cosine similarity between each pair of real and generated images and report the average across all samples $n$ as the $PSS_{element}$.
Scene-level Evaluation is designed to evaluate the consistency of the entire environmental scene (e.g., “forest,” “beach,” or “residential area”) between the generated landscape images and their corresponding ground truth images [35]. To infer scene categories, we employ a ResNet50 model [27], pre-trained on the Places365 dataset [83], to identify 365 predefined scene categories. For each soundscape $s$, we predict the top-$k$ scene labels and evaluate whether the generated image $\hat{l_{i}}$ aligns with the same scene category as its ground truth image $l_{i}$. The $PSS_{scene}$ is then calculated by its top-$k$ accuracy, where $k=1$ or $5$. $n$ is the total number of images, $P_{i}^{\text{k}}$ denotes the top-$k$ predicted category set for the $i$-th generated image, $T_{i}^{\text{k}}$ denotes the top-$k$ ground truth category set for the $i$-th real image.
Human Perception-level Evaluation is designed to evaluate whether the generated and real landscape images evoke similar human subjective perceptions and feelings, such as whether a place makes people feel “safe” or “lively” [82]. We utilize a DenseNet121 model [31], pre-trained on the MIT Place Pulse dataset [55], to measure six dimensions of human perceptions of environment, including “Safe,” “Beautiful,” “Depressing,” “Lively,” “Wealthy,” and “Boring”. For each image $i$, we compute the perception scores in six dimensions for the generated image $\mathrm{R}(\hat{l}_{i})$ and its corresponding ground truth $\mathrm{R}(l_{i})$. $PSS_{perception}$ is then defined as the average L1 distance across six dimensions, where $n$ is the number of the test set, $R$ is the Place Perception scores of each dimension.
By integrating metrics across element-, scene-, and human perception levels (Table 2), we provide a holistic geo-contextual evaluation framework. Informed by geographic domain knowledge and urban planning practices, we extend beyond measuring visual quality to assessing whether generated landscape images are contextually aligned with the environmental characteristics in input soundscapes.
6 Experiments
Figure 5: Visual comparison of landscape images generated by baseline models (CoDi, Sound2Scene, AudioToken Pre-trained/SD1/SD2, GlueGen) and our proposed SounDiT across SoundingSVI, SonicUrban, and two benchmark datasets. Ground truth landscape images are provided in the last column. Our proposed SounDiT captures significantly more geographic contexts than baselines.
6.1 Implementation Details
Model Architecture and Training. We adopt the pre-trained Variational Autoencoder (VAE) modules from Latent Stable Diffusion [65] as our landscape encoder–decoder pair, which were originally trained on the COCO dataset [45]. For soundscape conditioning, we employ the ImageBind-Huge audio encoder [22], pre-trained on two million AudioSet clips [21] covering a wide range of environmental soundscapes. We integrate a scene encoder based on the T5 architecture [62], which transforms high-level semantic scene prompts into latent embeddings to provide geographic semantics. Our SounDiT Block is initialized with the weights of PixArt-$\alpha$, a pre-trained text-to-image DiT model [10], and four experts are employed in TA-MoE Module. All components are trained using a learning rate of $1\times 10^{-4}$. Experiments are conducted on NVIDIA A100 and A6000 GPUs.
Benchmark Datasets. We curated two additional GeoS2L benchmark datasets to evaluate model generalizability. The first is a geo-contextual subset of VGGSound [9], comprising 16 classes with environmental semantics. It contains 6,713 soundscape–landscape pairs. The second combines the Landscape [41] and IntoTheWild [42] datasets, resulting in 1,889 soundscape-landscape pairs.
Baseline Models. We selected four state-of-the-art audio2image models, including CoDi [72], Sound2Scene [70], AudioToken [80], and GlueGen [61] as baseline models. For CoDi, we utilized its official pre-trained model. For AudioToken, we employed the pre-trained model without additional training, and trained it on our datasets using Stable Diffusion 1 and 2 as its backbone, respectively. The other two baselines are training from scratch based on their released code and training parameters.
Benchmark Evaluation Metrics. To evaluate the model performance, we adopted FID [30], AIS [30], and IIS [5], effective in previous audio-conditioned image generation studies [5, 71]. FID compares Gaussian distributions of features extracted from real and generated images, using a pre-trained Inception model. AIS measures audio-image semantic similarity using Wav2CLIP [78]. IIS evaluates image-image semantic similarity using CLIP-based embedding [29].
6.2 Main Results
[…]
General Metrics Place Similarity Score
[…]
Sound2Scene 78.137 0.586 0.623 0.467 0.374 0.735
[…]
AudioToken (SD1) 105.472 0.509 0.572 0.527 0.553 0.797
AudioToken (SD2) 149.473 0.504 0.559 0.249 0.311 0.845
GlueGen 74.250 0.500 0.566 0.400 0.461 0.776
SounDiT (Ours) 28.173 0.542 0.663 0.652 0.578 0.728
[…]
Sound2Scene 49.681 0.503 0.498 0.332 0.324 0.829
[…]
AudioToken (SD1) 95.210 0.505 0.503 0.419 0.484 0.979
AudioToken (SD2) 165.014 0.502 0.504 0.298 0.455 0.980
GlueGen 64.434 0.498 0.499 0.405 0.481 0.805
SounDiT (Ours) 33.246 0.517 0.592 0.465 0.495 0.788
[…]
Sound2Scene 140.244 0.617 0.741 0.479 0.640 0.769
[…]
AudioToken (SD1) 138.785 0.585 0.695 0.364 0.534 0.846
AudioToken (SD2) 118.536 0.637 0.740 0.512 0.724 0.855
GlueGen 195.625 0.584 0.645 0.362 0.582 0.815
SounDiT (Ours) 155.471 0.628 0.744 0.521 0.741 0.757
[…]
Sound2Scene 122.460 0.600 0.596 0.287 0.378 0.770
[…]
AudioToken (SD1) 163.540 0.532 0.555 0.207 0.311 0.789
AudioToken (SD2) 157.706 0.583 0.587 0.251 0.404 0.803
GlueGen 156.777 0.536 0.547 0.233 0.249 0.839
SounDiT (Ours) 113.111 0.575 0.588 0.336 0.392 0.767
Figure 5 presents landscape images generated by our proposed SounDiT and baseline models across SoundingSVI, Sonic Urban, and two benchmark datasets. The results generated by SounDiT show greater geo-contextual similarity to ground truth images, highlighting potential applications of SounDiT in geography and urban planning.
To evaluate model performance, we adopt general metrics (FID, AIS, IIS) in addition to our proposed geo-contextual PSS framework. As reported in Table 3, SounDiT consistently outperforms baselines on both SoundingSVI and SonicUrban datasets, with notable FID improvements (74$\rightarrow$28 and 49$\rightarrow$33), and achieves superior or competitive results across all metrics. SounDiT consistently ranks highest under PSS, indicating its ability to ensure geographic coherence in generated images. These results validate the effectiveness of PSS in measuring geo-contextual coherence beyond general metrics. The performance drop of SounDiT on the smaller IntoTheWild&Landscape dataset further emphasizes the crucial needs for large-scale training datasets such as SoundingSVI and SonicUrban.
User Study. In addition to these quantitative metrics, we further conducted a user study to assess whether the generated landscape images align with human perceptions of geographic contexts. Specifically, participants were asked to complete two matching tasks: (1) selecting the generated image that best aligned with the given soundscape; and (2) identifying which generated landscape image most closely matched a ground-truth landscape. The average matching accuracy is 86.13%, indicating a strong perceptual alignment between soundscapes and their corresponding generated images. This demonstrates the effectiveness of SounDiT for GeoS2L.
6.3 Ablations
Table 4: Ablation results. We assess the contribution of key components to the overall performance of SounDiT.
[…]
TA-MoE 28.173 0.652
MHCA 34.108 0.602
(a) SounDiT Block Design
[…]
w/ RAG 0.652 0.578
w/o RAG 0.479 0.514
(b) Geo-contextual RAG
We conducted the ablation study on SoundingSVI. First, we replace the TA-PMoE module with a standard multi-head cross-attention layer, forming SounDiT-MHCA. Second, we disable the RAG module in SounDiT-TA-MoE, yielding SounDiT-TA-MoE w/o RAG. As reported in Table 4(a), SounDiT-TA-MoE achieves a lower FID and a higher $PSS_{element}$, indicating that the expert routing (TA-MoE) is essential for modeling multi-level soundscape and visual features, thereby preserving visual quality and geographic coherence in the generated images. Table 4(b) further shows that incorporating geo-contextual RAG inference improves both element- and scene-level PSS of generated landscape images, highlighting the benefits of integrating geographic contexts during inference.
7 Concluding Remarks
In this work, we present Geo-Contextual Soundscape-to-Landscape (GeoS2L), a practically-informed novel generation problem that synthesizes realistic landscape images from environmental soundscapes. We explicitly integrate geographic domain knowledge into multimodal generative modeling. To support GeoS2L, we construct SoundingSVI and SonicUrban, two large-scale, multi-modal, geo-contextual datasets, capturing diverse sounding and visual environments. We propose SounDiT, a novel diffusion-based framework that captures geographic semantics embedded in sounding environments. We further design a novel geo-contextual evaluation framework, the Place Similarity Score (PSS), which assesses geographic and environmental consistency. Extensive experiments demonstrate that SounDiT consistently outperforms existing baselines and can generate realistic landscape images. The proposed geo-contextual datasets, model, and evaluation framework effectively integrate geographic knowledge into generative modeling, establishing a solid benchmark for the GeoS2L task.
Limitations. While our work introduces a novel problem, GeoS2L, several limitations remain. First, the SoundingSVI dataset relies on geotagged audio from Aporee, which is a crowd-sourced platform that may contain spatial inaccuracies. Future efforts may focus on refining spatial alignment and improving dataset quality through more robust data processing procedures. Second, the current PSS evaluation metrics only consider three practically-informed levels. Future studies could integrate more geographic domain-specific knowledge to support geo-contextually coherent image generation.